OpenStudy (anonymous):

a. Given the function y1 = x2 + 4x-12 Calculate -if applicable ─ intersections with the axes, minimum and maximum and the domain of the function.

OpenStudy (anonymous):

\[x ^{2}\]

OpenStudy (anonymous):

not x2

OpenStudy (phi):

The domain are all "allowed" x values usually every x is allowed unless it causes trouble (for example divide by zero, or square root of a negative number). For this equation, no x causes trouble, so the domain is -infty to + infty, which you can write as \( (-\infty , \infty) \)

OpenStudy (phi):

For intersections with the axes for the y-axis, this happens when x=0. So put 0 in for x in the equation and simplify to get the y value where the curve intersects the y-axis

OpenStudy (anonymous):

ahh so far i understand it now

OpenStudy (phi):

what do you get for the intersection with the y-axis?

OpenStudy (anonymous):

y=(0,-12)

OpenStudy (phi):

intersection with the x-axis happens when y=0. you have to solve for x : \[ x^2 + 4x-12=0 \] this quadratic factors

OpenStudy (phi):

can you factor it ?

OpenStudy (phi):

for the max value, we can focus on the "highest order term", i.e. the x^2 term as x gets large, x^2 is even larger. And because x can "go to infinity" , so does x^2 that means the max value is +infinity (or unbounded)

OpenStudy (phi):

for the min value, this curve is a parabola with a "smile" \( \cup\) shape its smallest value is at its vertex, and the x value of the vertex is at \[ x_{vertex}= \frac{-b}{2a} \] where to find a and b, match your equation with \[ y= a x^2 + bx +c \]

OpenStudy (anonymous):

I do not know if i can factor it... i do not know how to solve that one.

OpenStudy (phi):

Something tells me you already had that lesson. But you do this: look at the last number, the -12 list all pairs of numbers that when you multiply them, you get 12 they are: 1,12 2,6 3,4 Next, the - in front of the 12 means the factors will have different signs (one + the other -) look at the number in front of the x, i.e. +4 the + means the larger number in each pair is positive. so put in the signs: -1+12 -2+4 -3+4 now add each pair. if any pair adds up to +4 , those are the factors

OpenStudy (phi):

**the middle pair should be -2+6

OpenStudy (anonymous):

the middle pair adds up to +4

OpenStudy (phi):

-2+6 = 4 so -2 and +6 go into (x-2)(x+6) if we multiply that out, we get x^2 +4x -12, which is what we started with so the problem \[ x^2 +4x -12=0 \] can be written as \[ (x-2)(x+6)=0 \] now it is easy to see which x values make the left side zero. any ideas?

OpenStudy (anonymous):

2 and 6?

OpenStudy (phi):

let's test 2: (2-2)(2+6) 0*8 = 0 yes that worked now 6: (6-2)*(6+6) 4*12= 48. no, 6 does not work

OpenStudy (anonymous):

-6?

OpenStudy (phi):

-6 is much better , because -6+6 is 0 so the x-intercepts are at x=-6 and x=2

OpenStudy (anonymous):

ok thanks :)

OpenStudy (phi):

Can you find the min value? You have to first find the x value of the vertex (see up above)

OpenStudy (anonymous):

is it -24? i did \[-4\div2 = -2\] then used -2 in the formula

OpenStudy (phi):

yes, x=-2 y= x^2 + 4x-12 y= (-2)*(-2) + 4*-2 -12 = 4 -8 -12 = -4-12 = -16

OpenStudy (phi):

it looks like you did (-2)^2 and got -4, so try to figure out why you did the wrong. -2 * -2 is +4

OpenStudy (anonymous):

ahh yes I did not put -2 , sloppy error

OpenStudy (anonymous):

thanks!